# Assessing the Effects of Monetary Shocks on Macroeconomic Stars: A SMUC-IV Framework

### v1

@@ 2.2 Identification via External Instrument
In this section, we first discuss how the external instruments can be incorporated into our modeling framework for identification, and then describe our proposed SMUC-IV as an augmented strucutrual vector autoregression, a form which is used later for Bayesian estimation. In this paper, we consider the case in which one external instrument is used to identify one structural shock, namely, the monetary policy shock of interest. For the case of using multiple external instruments to identify multiple structural shocks, we refer readers to Arias et al., (2021), Braun and Brüggemann, (2023), and Hou, (2024) for more details.
@@
These two conditions are central to understanding how the external instrument can be used to identify the structural shock of interest $\epsilon_{m,t}$. Specifically, it conveys identifying information by distinguishing $\epsilon_{m,t}$ from the other shocks in $\bm{\epsilon}_{-m,t}$ through differences in their covariance structures with the the external instrument $m_{t}$.
@@
To summarize, our SMUC-IV is specified as equations (9), (10), and (11), subject to the zero restrictions given in (12). The external instrument equation (11) with the zero restrictions in (12) gives $m_{t}=\beta\epsilon_{m,t}+\alpha v_{t}$, which is consistent with the specification used in Caldara and Herbst, (2019).
@@
As indicated by recent studies (Arias et al., , 2021; Braun and Brüggemann, , 2023; Hou, , 2024), the use of external instruments can only achieve set-identification. For instance, in our case, it can be shown that the second to last column of $\widetilde{\mathbf{B}}$ that embeds the impact responses of monetary policy shocks, i.e., the last column of $\mathbf{B}$, which can only be identified up to sign changes. A common solution to this set-identification issue is to impose sign restrictions on the impulse responses, which are informed by economic theory. In this paper, rather than imposing dogmatic restrictions on response directions, we identify monetary policy shocks using an informative prior that accommodates estimation uncertainty. We will discuss this more in Section 3.1.
@@ 3.1 Priors
This section describes the prior distributions assigned to the model parameters. Let $\bar{\bm{\tau}}_{0}=(g_{-1}^{*},g_{0}^{*},\pi_{0}^{*},r_{0}^{*})^{\prime}$ denote the vector of the initial state parameters. Let $\bm{\Phi}_{l,i,j}$ represent the $(i,j)$ element of the autoregressive coefficient matrix $\bm{\Phi}_{l}$ for $i,j=1,2,3$, $l=1,\ldots,p$, and let $\mathbf{B}_{i,j}$ denote the $(i,j)$ element of the contemporaneous impact matrix $\mathbf{B}$ for $i,j=1,\ldots,6$. We assume the following independent priors:
@@
This specification reflects the prior belief that the coefficients on more distant lags are less important than those on recent lags, and are therefore shrunk more strongly toward zero. Following standard practice, the scale parameter $\sigma^{2}_{i}$ is set equal to the residual variance of an AR($p$) model for its corresponding variable $i$. The hyperparameters $\kappa_{1}$ and $\kappa_{2}$ control the shrinkage strength for the own-lag and cross-variable-lag coefficients, respectively. Empirical evidence from recent studies, such as Cross et al., (2020) and Chan, (2021), indicates that allowing the hyperparameters of shrinkage prior to be parameters to be estimated can substantially improves both forecasting accuracy and model fit. Therefore, rather than fixing these hyperparameters at predetermined values, we treat $\kappa_{1}$ and $\kappa_{2}$ as unknown parameters and assign them the following prior distributions:
@@
The prior we consider here provides additional information for identifying the contemporaneous response matrix $\mathbf{B}$. Recall that the monetary policy shock $\epsilon_{m,t}$ is ordered last in $\bm{\epsilon}_{t}$. That means, the last column of $\mathbf{B}$ is the impact responses of the trend and cycle components to a monetary policy shock. Our prior reflects a belief that a one-standard-deviation increase in monetary policy shock $\epsilon_{m,t}$ results in a $100$ basis point increase in the cycle of interest rate $c_{r,t}$, while having no effect on $\Delta g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$, on impact. By the definition of $c_{r,t}$ in (3), the impact response of $c_{r,t}$ to a monetary policy shock is entirely attributable to the increase in $r_{t}$. This identifying information from our prior is similar to conventional studies on identifying monetary policy shock by normalize the magnitude of the interest rate response to a positive value. [^4] For instance, Bauer and Swanson, (2023) normalize the interest rate response to a monetary policy shock by $25$ basis points, while Miranda-Agrippino and Ricco, (2021) use a normalization of $100$ basis points. Unlike these approaches, we do not fix the size of the response of interest rate to a monetary policy shock, instead, we impose a relatively tight prior that centers the response at $100$ basis points. We also highlight that since we use an informative prior that centers the impact responses of $\Delta g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$ to a monetary policy shock at zeros, any nonzero effects found in our empirical study must be supported by the data.
For the parameters in the external instrument equation, we set $V_{\beta}=1$, $V_{\alpha}=1$, $\alpha_{0}=0$, and $\beta_{0}=0.5\times\sigma_{m}$, where $\sigma_{m}$ denotes the standard deviation of the external instrument. This is comparable to the setting in Caldara and Herbst, (2019). For the initial state parameters $\bar{\bm{\tau}}_{0}$, we use an uninformative prior by setting $\mathbf{V}_{\bar{\tau}_{00}}=100\mathbf{I}_{4}$ and $\bar{\bm{\tau}}_{00}=(g_{1},g_{1},\pi_{1},r_{1})^{\prime}$ where $g_{1}$, $\pi_{1}$ and $r_{1}$ are the first observations of log real GDP, inflation, and the interest rate in the sample period, respectively.
@@ 3.2 Posterior Sampler
We now discuss the estimation of our model with the prior described in the previous section. To set the stage, let $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, where $\mathbf{y}_{t}=(g_{t},\pi_{t},r_{t},m_{t})^{\prime}$, denote the vector of observed data, $\bm{\Phi}=(\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})$ denote the collection of the autoregressive coefficient matrices and $\bm{\tau}=(\bar{\bm{\tau}}_{0}^{\prime},\bm{\tau}_{1}^{\prime},\ldots,\bm{\tau}_{T}^{\prime})^{\prime}$ denote a vector of state parameters. The joint posterior distribution can be simulated by sequentially sampling from the following conditional distributions:
@@
In Step 1, we sample the state vector $\bm{\tau}$ using the precision sampling method (Chan and Jeliazkov, , 2009; McCausland et al., , 2011; Rue, , 2001) instead of traditional Kalman filter techniques. The precision-based sampling method has gain increasing popularity for estimating state space models due to its computational efficiency and straightforward implementation. One complication under our framework, however, involves determining the conditional posterior distribution for $\bm{\tau}$. The conventional approach applies Bayes’ rule, which requires separately deriving both the conditional likelihood and the prior density of the state parameters. However, when the state and measurement equations are correlated, the derivation of the conditional posterior of the state parameters becomes more cumbersome (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Leiva-Leon and Uzeda, , 2023).
In this paper, we introduce a novel and direct approach to derive the conditional posterior of $\bm{\tau}$ that circumvents the use of Bayes’ rule. The key idea is to first obtain the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$, which will be shown later to be a Gaussian distribution, and then exploit the standard Gaussian conditioning properties to obtain the conditional posterior of $\bm{\tau}$. Specifically, by stacking equation (13) over $t=1,\ldots,T$, we obtain
@@
$\mathbf{H}_{1}=\begin{pmatrix}\mathbf{I}_{7}&\mathbf{0}_{7}&\mathbf{0}_{7}&\cdots&\cdots&\cdots&\cdots&\mathbf{0}_{7}\\ -\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{p}&\ddots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\vdots\\ \mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\mathbf{0}_{7}\\ \mathbf{0}_{7}&\cdots&\mathbf{0}_{7}&-\widetilde{\mathbf{A}}_{p}&\cdots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}\end{pmatrix},\quad\bm{\Xi}=\begin{pmatrix}\bm{\Xi}_{1}\\ \mathbf{0}_{4\times 4}\\ \bm{\Xi}_{2}\\ \mathbf{0}_{4\times 4}\\ \mathbf{0}_{7\times 4}\\ \vdots\\ \mathbf{0}_{7\times 4}\end{pmatrix}$
with
$\bm{\Xi}_{1}=\begin{pmatrix}-1&2&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix},\quad\bm{\Xi}_{2}=\begin{pmatrix}0&-1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$.
@@
where $\widehat{\bm{\tau}}=\bm{\mu}_{\tau}-\mathbf{K}_{\tau}^{-1}\mathbf{K}_{\tau,y}(\mathbf{y}-\bm{\mu}_{y})$. It can be check that the precision matrix $\mathbf{K}_{\tau}$ is a banded matrix. Therefore, one can efficiently sample from $\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ using the precision-based approach of Chan and Jeliazkov, (2009).
For sampling $(\mathbf{B},\beta,\alpha)$ in Step 2, we first note that these parameters correspond to the nonzero elements in the contemporaneous impact matrix $\widetilde{\mathbf{B}}$ of the augmented structural autoregression representation in (13). Consequently, sampling $(\mathbf{B},\beta,\alpha)$ is equivalent to sampling $\widetilde{\mathbf{B}}$ with the zero restrictions in (15) and the sign restriction on $\alpha$. Since these restrictions can be expressed as linear equality and inequality restrictions on $\widetilde{\mathbf{B}}$, we sample the nonzero parameters using the efficient sampler developed by Hou, (2024). We refer readers to this paper for more implementation details.
@@ 4 Marginal Likelihood Estimation
Marginal likelihood is the standard criterion for Bayesian model comparison. In this section, we present an approach for estimating the marginal likelihood of our proposed SMUC-IV and its restricted versions. Our method builds upon the modified harmonic mean estimator proposed by Gelfand and Dey, (1994), integrating it with the conditional Monte Carlo method of Chan, (2023) to enhance estimation accuracy. We begin with an overview of the modified harmonic mean estimator and then outline our conditional Monte Carlo improved estimator for the SMUC-IV. Technical details are relegated to Appendix B.
@@ 4.1 Modified Harmonic Mean Estimator
While the modified harmonic mean estimator described above is simulation-consistent and straightforward to implement, it might perform poorly when the model parameter $\bm{\theta}$ is high-dimensional. To improve estimation accuracy, we employ the conditional Monte Carlo method of Chan, (2023). The key idea is to first analytically integrate out as many parameters in $\bm{\theta}$ as possible, and then construct a modified harmonic mean estimator using the resulting unconditional likelihood and prior on the remaining lower-dimensional parameters.

### v2

@@ 2.2 Identification via External Instrument
In this section, we first discuss how the external instrument can be incorporated into our modelling framework for identification, and then describe our proposed SMUC-IV as an augmented structural vector autoregression, a form which is used later for Bayesian estimation. In this paper, we consider the case in which one external instrument is used to identify one structural shock, namely, the monetary policy shock of interest. For the case of using multiple external instruments to identify multiple structural shocks, we refer readers to arias2021inference, braun2023identification, and hou2024large for more details.
@@
These two conditions are central to understanding how the external instrument can be used to identify the structural shock of interest $\epsilon_{m,t}$. Specifically, it conveys identifying information by distinguishing $\epsilon_{m,t}$ from the other shocks in $\bm{\epsilon}_{-m,t}$ through differences in their covariance structures with the external instrument $m_{t}$.
@@
To summarize, our SMUC-IV is specified as equations (9), (10), and (11), subject to the zero restrictions given in (12). The external instrument equation (11) with the zero restrictions in (12) gives $m_{t}=\beta\epsilon_{m,t}+\alpha v_{t}$, which is consistent with the specification used in caldara2019monetary.
@@
As indicated by recent studies (arias2021inference; braun2023identification; hou2024large), the use of external instruments can only achieve set-identification. For instance, in our case, it can be shown that the second-to-last column of $\widetilde{\mathbf{B}}$, which embeds the impact responses to the monetary policy shock and corresponds to the last column of $\mathbf{B}$, can only be identified up to a sign change. A common solution to this set-identification issue is to impose sign restrictions on the impulse responses, which are informed by economic theory. In this paper, rather than imposing dogmatic restrictions on response directions, we identify monetary policy shocks using an informative prior that accommodates estimation uncertainty. We will discuss the prior in more detail in Section 3.1.
@@ 3.1 Priors
This section describes the prior distributions assigned to the model parameters. Let $\bar{\bm{\tau}}_{0}=(g_{0}^{*},\pi_{0}^{*},r_{0}^{*})^{\prime}$ denote the vector of the initial state parameters, $\bm{\Phi}_{l,i,j}$ the $(i,j)$th element of the autoregressive coefficient matrix $\bm{\Phi}_{l}$ for $i,j=1,2,3$, $l=1,\ldots,p$, and $\mathbf{B}_{i,j}$ the $(i,j)$th element of the contemporaneous impact matrix $\mathbf{B}$ for $i,j=1,\ldots,6$. We assume the following independent priors:
@@
This specification reflects the prior belief that the coefficients on more distant lags are less important than those on recent lags, and are therefore shrunk more strongly toward zero. Following standard practice, the scale parameter $\sigma^{2}_{i}$ is set equal to the residual variance of an AR($p$) model for its corresponding variable $i$. The hyperparameters $\kappa_{1}$ and $\kappa_{2}$ control the shrinkage strength for the own-lag and cross-variable-lag coefficients, respectively. Empirical evidence from recent studies, such as cross2020macroeconomic and chan2021minnesota, indicates that allowing the hyperparameters of shrinkage prior to be parameters to be estimated can substantially improves both forecasting accuracy and model fit. Therefore, rather than fixing these hyperparameters at predetermined values, we treat $\kappa_{1}$ and $\kappa_{2}$ as unknown parameters and assign them the following prior distributions:
@@
The prior we consider here provides additional information for identifying the contemporaneous response matrix $\mathbf{B}$. Recall that the monetary policy shock $\epsilon_{m,t}$ is ordered last in $\bm{\epsilon}_{t}$. That means, the last column of $\mathbf{B}$ is the impact responses of the trend and cycle components to a monetary policy shock. Our prior reflects a belief that a one-standard-deviation increase in monetary policy shock $\epsilon_{m,t}$ results in a $100$ basis point increase in the cycle of nominal interest rate $c_{i,t}$, while having no effect on $g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$, on impact. By the definition of $c_{i,t}$ in (3), the impact response of $c_{i,t}$ to a monetary policy shock is entirely attributable to the increase in $i_{t}$. This identifying information from our prior is similar to conventional studies on identifying monetary policy shock by normalize the magnitude of the nominal interest rate response to a positive value. [^4] For instance, bauer2023reassessment normalize the nominal interest rate response to a monetary policy shock by $25$ basis points, while miranda2021transmission use a normalization of $100$ basis points. Unlike these approaches, we do not fix the size of the response of nominal interest rate to a monetary policy shock, instead, we impose a relatively tight prior that centers the response at $100$ basis points. We also highlight that since we use an informative prior that centers the impact responses of $g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$ to a monetary policy shock at zeros, any nonzero effects found in our empirical study must be supported by the data.
For the parameters in the external instrument equation, we set $V_{\beta}=1$, $V_{\alpha}=1$, $\alpha_{0}=0$, and $\beta_{0}=0.5\times\sigma_{m}$, where $\sigma_{m}$ denotes the standard deviation of the external instrument. This is comparable to the setting in caldara2019monetary. For the initial state parameters $\bar{\bm{\tau}}_{0}$, we use an uninformative prior by setting $\mathbf{V}_{\bar{\tau}_{00}}=100\mathbf{I}_{3}$ and $\bar{\bm{\tau}}_{00}=(g_{1},\pi_{1},i_{1})^{\prime}$ where $g_{1}$, $\pi_{1}$ and $i_{1}$ are the first observations of log real GDP, inflation, and the nominal interest rate in the sample period, respectively.
@@ 3.2 Posterior Sampler
We now discuss the estimation of our model with the prior described in the previous section. To set the stage, let $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, where $\mathbf{y}_{t}=(g_{t},\pi_{t},i_{t},m_{t})^{\prime}$, denote the vector of observed data, $\bm{\Phi}=(\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})$ denote the collection of the autoregressive coefficient matrices and $\bm{\tau}=(\bar{\bm{\tau}}_{0}^{\prime},\bm{\tau}_{1}^{\prime},\ldots,\bm{\tau}_{T}^{\prime})^{\prime}$ denote a vector of state parameters. The joint posterior distribution can be simulated by sequentially sampling from the following conditional distributions:
@@
In Step 1, we sample the state vector $\bm{\tau}$ using the precision sampling method (chan2009efficient; mccausland2011simulation; rue2001fast) instead of traditional Kalman filter techniques. The precision-based sampling method has gained increasing popularity for estimating state space models due to its computational efficiency and straightforward implementation. One complication under our framework, however, involves determining the conditional posterior distribution for $\bm{\tau}$. The conventional approach requires separately deriving both the conditional likelihood and the prior density of the state parameters. However, when the state and measurement equations are correlated, the derivation of the conditional posterior of the state parameters becomes more cumbersome (grant2017bayesian; grant2017reconciling; leiva2023endogenous).
In this paper, we introduce a novel and direct approach to derive the conditional posterior of $\bm{\tau}$. The key idea is to first obtain the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$, which will be shown later to be a Gaussian distribution, and then exploit the standard Gaussian conditioning properties to obtain the conditional posterior of $\bm{\tau}$. Specifically, by stacking equation (13) over $t=1,\ldots,T$, we obtain
@@
$\mathbf{H}_{1}=\begin{pmatrix}\mathbf{I}_{7}&\mathbf{0}_{7}&\mathbf{0}_{7}&\cdots&\cdots&\cdots&\cdots&\mathbf{0}_{7}\\ -\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{p}&\ddots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\vdots\\ \mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\mathbf{0}_{7}\\ \mathbf{0}_{7}&\cdots&\mathbf{0}_{7}&-\widetilde{\mathbf{A}}_{p}&\cdots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}\end{pmatrix},\quad\bm{\Xi}=\begin{pmatrix}\mathbf{I}_{3}\\ \mathbf{0}_{4\times 3}\\ \mathbf{0}_{7\times 3}\\ \vdots\\ \vdots\\ \mathbf{0}_{7\times 3}\end{pmatrix}$
@@
where $\widehat{\bm{\tau}}=\bm{\mu}_{\tau}-\mathbf{K}_{\tau}^{-1}\mathbf{K}_{\tau,y}(\mathbf{y}-\bm{\mu}_{y})$. Note that the $(3T+3)\times(3T+3)$ precision matrix $\mathbf{K}_{\tau}$ is sparse and banded, as illustrated in Figure 1, which visualizes its structure with nonzero and zero elements shown in blue and white, respectively. [^5] Here the dimension of $\mathbf{K}_{\tau}$ is $432\times 432$, which matches the dimension used in our empirical application in Section 5. This banded structure enables efficient sampling from $\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ using the precision-based approach of chan2009efficient. The procedure is as follows: first, we compute the lower triangular Cholesky factor $\mathbf{L}$ of $\mathbf{K}_{\tau}$ such that $\mathbf{K}_{\tau}=\mathbf{L}\mathbf{L}^{\prime}$, which is computationally efficient due to the banded structure of $\mathbf{K}_{\tau}$. A draw $\bm{\tau}\sim\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ is then obtained by solving:
$\bm{\tau}=\bm{\mu}_{\tau}-\mathbf{L}^{{}^{\prime}-1}\left(\mathbf{L}^{-1}\mathbf{K}_{\tau,y}(\mathbf{y}-\bm{\mu}_{y})+\mathbf{f}\right),\quad\mathbf{f}\sim\mathcal{N}\left(\mathbf{0}_{3(T+1)\times 1},\mathbf{I}_{3(T+1)}\right).$
Since $\mathbf{L}$ is a banded lower triangular matrix, we can solve the required linear systems rapidly via forward and backward substitution, with $\mathcal{O}(T)$ complexity. This efficiently avoids directly computing $\mathbf{K}_{\tau}^{-1}$, which has $\mathcal{O}(T^{3})$ complexity.
Figure 1: The sparsity pattern of the precision matrix $\mathbf{K}_{\tau}$.
For sampling $(\mathbf{B},\beta,\alpha)$ in Step 2, we first note that these parameters correspond to the nonzero elements in the contemporaneous impact matrix $\widetilde{\mathbf{B}}$ of the augmented structural autoregression representation in (13). Consequently, sampling $(\mathbf{B},\beta,\alpha)$ is equivalent to sampling $\widetilde{\mathbf{B}}$ with the zero restrictions in (15) and the sign restriction on $\alpha$. Since these restrictions can be expressed as linear equality and inequality restrictions on $\widetilde{\mathbf{B}}$, we sample the nonzero parameters using the efficient sampler developed by hou2024large. We refer readers to this paper for more implementation details.
@@ 4 Marginal Likelihood Estimation
Marginal likelihood is the standard criterion for Bayesian model comparison. In this section, we present an approach for estimating the marginal likelihood of our proposed SMUC-IV and its restricted versions. Our method builds upon the modified harmonic mean estimator proposed by gelfand1994bayesian, integrating it with the conditional Monte Carlo method of chan2023comparing to enhance estimation accuracy. We begin with an overview of the modified harmonic mean estimator and then outline our conditional Monte Carlo improved estimator for the SMUC-IV. Technical details are relegated to Appendix B.
@@ 4.1 Modified Harmonic Mean Estimator
While the modified harmonic mean estimator described above is simulation-consistent and straightforward to implement, it might perform poorly when the model parameter $\bm{\theta}$ is high-dimensional. To improve estimation accuracy, we employ the conditional Monte Carlo method of chan2023comparing. The key idea is to first analytically integrate out as many parameters in $\bm{\theta}$ as possible, and then construct a modified harmonic mean estimator using the resulting unconditional likelihood and prior on the remaining lower-dimensional parameters.